please help me compute the final step to this question, and if possible be as detailed as you can, I am unsure what step I am messing up.Suppose 6% of students are veterans and 126 students are involved in sports. How unusual would it be to have no more than 12
Answer #1

Solution

\sigma\hat p =  \sqrt[p ( 1 - p ) / n] = 0.021157009

P( \hat p \le 0.095238) =

= P[( \hat p - \mu \hat p ) / \sigma \hat p\le (0.095238 - 0.06) / 0.021157009]

= P(z\le1.666)

= 0.9521